Please help with my problem #1d. Solve the equation:
4 - (6 + x) = 13
Thank you.

Answer by longjonsilver(2297)

4 - (6 + x) = 13
4 - 6 - x = 13
-2 - x = 13
-x = 15
x = -15

jon.